# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes

### v1

@@ Appendix C Additional results
covid (n=40) 77.93 71.57 75.23 75.36 normal-$\uparrow$ 71.55 73.37 $\uparrow$ 66.17 70.85 $\uparrow$ 72.21 72.68 $\uparrow$ 56.03 70.80 $\uparrow$ 67.39 71.82 $\uparrow$ 70.02 69.96 61.19 69.77 $\uparrow$
cutract (n=40) 73.83 69.50 71.26 73.57 normal-$\uparrow$ 69.66 71.34 $\uparrow$ 62.52 68.69 $\uparrow$ 64.50 68.05 $\uparrow$ 51.23 68.23 $\uparrow$ 60.74 68.91 $\uparrow$ 69.42 69.19 59.34 67.80 $\uparrow$
maggic (n=40) 71.45 58.76 60.27 61.88 normal-$\uparrow$ 57.87 60.29 $\uparrow$ 55.87 58.47 $\uparrow$ 55.38 58.03 $\uparrow$ 48.50 57.02 $\uparrow$ 54.53 59.01 $\uparrow$ 58.12 58.29 $\uparrow$ 56.19 58.33 $\uparrow$
seer (n=40) 91.79 84.51 86.63 87.40 normal-$\uparrow$ 86.27 86.80 $\uparrow$ 82.48 86.55 $\uparrow$ 81.62 81.58 52.80 83.03 $\uparrow$ 80.04 85.84 $\uparrow$ 82.53 83.62 $\uparrow$ 79.35 85.81 $\uparrow$
compas (n=40) 69.26 62.90 69.24 68.90 61.74 63.21 $\uparrow$ 56.63 62.55 $\uparrow$ 59.07 59.38 $\uparrow$ 58.01 64.95 $\uparrow$ 58.44 64.44 $\uparrow$ 61.81 61.89 $\uparrow$ 56.60 62.26 $\uparrow$
adult (n=40) 88.88 82.16 48.51 76.59 $\uparrow$ 45.92 76.92 $\uparrow$ 73.38 82.53 $\uparrow$ 74.03 79.84 $\uparrow$ 66.60 80.98 $\uparrow$ 69.47 81.56 $\uparrow$ 80.01 80.51 $\uparrow$ 74.20 82.53 normal-$\uparrow$
drug (n=40) 80.63 72.72 77.03 78.85 normal-$\uparrow$ 76.37 76.48 $\uparrow$ 70.99 74.48 $\uparrow$ 74.65 76.53 $\uparrow$ 64.52 75.09 $\uparrow$ 62.58 73.14 $\uparrow$ 71.80 72.15 $\uparrow$ 67.61 74.04 $\uparrow$
covid (n=100) 77.70 73.16 75.55 76.56 normal-$\uparrow$ 72.69 74.36 $\uparrow$ 70.54 73.99 $\uparrow$ 72.99 74.50 $\uparrow$ 63.40 73.98 $\uparrow$ 63.50 73.79 $\uparrow$ 72.64 73.35 $\uparrow$ 70.45 73.76 $\uparrow$
cutract (n=100) 73.04 70.49 72.29 74.01 $\uparrow$ 71.51 73.48 $\uparrow$ 68.64 73.63 $\uparrow$ 71.32 74.02 normal-$\uparrow$ 54.74 69.69 $\uparrow$ 66.97 72.98 $\uparrow$ 67.71 69.42 $\uparrow$ 66.44 72.32 $\uparrow$
maggic (n=100) 71.42 60.55 60.04 62.57 normal-$\uparrow$ 56.98 60.49 $\uparrow$ 58.21 60.42 $\uparrow$ 58.11 59.42 $\uparrow$ 49.71 58.77 $\uparrow$ 55.36 60.58 $\uparrow$ 60.02 60.15 $\uparrow$ 56.47 60.28 $\uparrow$
seer (n=100) 91.81 85.52 86.81 87.64 normal-$\uparrow$ 86.39 87.21 $\uparrow$ 83.74 87.29 $\uparrow$ 84.66 86.85 $\uparrow$ 68.20 85.49 $\uparrow$ 82.37 86.80 $\uparrow$ 83.18 83.87 $\uparrow$ 80.45 86.49 $\uparrow$
compas (n=100) 68.88 62.71 69.37 69.37 59.83 62.80 $\uparrow$ 60.56 64.90 $\uparrow$ 58.20 62.00 $\uparrow$ 58.51 64.60 $\uparrow$ 59.76 64.68 $\uparrow$ 61.33 62.03 $\uparrow$ 60.37 63.42 $\uparrow$
adult (n=100) 88.71 82.78 44.60 76.75 $\uparrow$ 46.34 76.83 $\uparrow$ 74.97 82.89 $\uparrow$ 77.38 81.42 $\uparrow$ 78.84 83.96 normal-$\uparrow$ 68.31 82.87 $\uparrow$ 82.58 83.24 $\uparrow$ 76.15 82.47 $\uparrow$
drug (n=100) 81.44 75.49 79.12 79.81 normal-$\uparrow$ 72.52 74.78 $\uparrow$ 68.89 75.91 $\uparrow$ 76.88 78.33 $\uparrow$ 66.38 77.02 $\uparrow$ 62.27 75.13 $\uparrow$ 74.40 74.49 $\uparrow$ 71.02 76.85 $\uparrow$
covid (n=200) 77.56 73.06 75.12 76.64 normal-$\uparrow$ 71.74 74.66 $\uparrow$ 72.57 75.50 $\uparrow$ 73.56 75.88 $\uparrow$ 66.02 74.51 $\uparrow$ 63.29 74.62 $\uparrow$ 73.34 73.89 $\uparrow$ 68.14 74.21 $\uparrow$
cutract (n=200) 73.68 71.68 73.43 75.14 $\uparrow$ 71.76 74.31 $\uparrow$ 69.35 73.80 $\uparrow$ 72.76 75.33 normal-$\uparrow$ 65.83 73.43 $\uparrow$ 68.05 74.27 $\uparrow$ 70.06 71.35 $\uparrow$ 66.65 73.34 $\uparrow$
maggic (n=200) 71.41 62.40 59.46 63.15 normal-$\uparrow$ 58.05 61.68 $\uparrow$ 59.04 62.58 $\uparrow$ 57.63 59.23 $\uparrow$ 50.63 60.09 $\uparrow$ 55.55 61.36 $\uparrow$ 61.73 61.88 $\uparrow$ 57.09 61.15 $\uparrow$
seer (n=200) 91.83 85.87 86.75 88.08 $\uparrow$ 85.74 86.97 $\uparrow$ 83.97 87.64 $\uparrow$ 86.21 88.46 normal-$\uparrow$ 83.21 87.65 $\uparrow$ 76.91 87.16 $\uparrow$ 84.70 85.59 $\uparrow$ 81.45 87.46 $\uparrow$
compas (n=200) 69.04 63.14 68.80 69.54 normal-$\uparrow$ 61.62 64.59 $\uparrow$ 60.65 65.67 $\uparrow$ 60.97 65.12 $\uparrow$ 60.66 65.67 $\uparrow$ 56.49 64.34 $\uparrow$ 61.68 62.93 $\uparrow$ 60.32 64.18 $\uparrow$
adult (n=200) 88.86 84.12 40.38 77.39 $\uparrow$ 47.40 79.40 $\uparrow$ 80.00 84.61 $\uparrow$ 83.73 85.73 $\uparrow$ 81.92 85.76 normal-$\uparrow$ 74.90 85.13 $\uparrow$ 84.78 84.45 75.54 84.39 $\uparrow$
drug (n=200) 80.13 78.41 78.25 79.09 $\uparrow$ 73.31 76.76 $\uparrow$ 72.82 79.00 $\uparrow$ 72.62 74.89 $\uparrow$ 74.30 79.75 normal-$\uparrow$ 63.91 77.60 $\uparrow$ 77.64 77.92 $\uparrow$ 70.16 77.89 $\uparrow$
@@
covid (n=20) 63.46 61.92 67.52 67.67 normal-$\uparrow$ 61.07 64.89 $\uparrow$ 52.94 56.62 $\uparrow$ 58.95 59.19 $\uparrow$ 51.83 58.73 $\uparrow$ 54.37 61.53 $\uparrow$ 59.56 59.01 56.46 59.49 $\uparrow$
cutract (n=20) 62.49 64.62 62.40 66.03 normal-$\uparrow$ 62.10 65.15 $\uparrow$ 55.67 61.73 $\uparrow$ 57.91 58.97 $\uparrow$ 51.89 62.76 $\uparrow$ 57.12 64.85 $\uparrow$ 62.85 62.38 60.25 62.34 $\uparrow$
maggic (n=20) 56.53 54.80 54.67 55.03 normal-$\uparrow$ 52.54 53.54 $\uparrow$ 50.90 52.08 $\uparrow$ 52.39 52.70 $\uparrow$ 49.70 53.08 $\uparrow$ 51.52 53.23 $\uparrow$ 52.55 52.44 50.91 51.52 $\uparrow$
seer (n=20) 78.86 71.09 76.17 77.17 normal-$\uparrow$ 73.92 75.58 $\uparrow$ 64.18 70.69 $\uparrow$ 74.78 74.72 46.60 68.79 $\uparrow$ 65.77 72.25 $\uparrow$ 70.54 71.19 $\uparrow$ 64.47 70.11 $\uparrow$
compas (n=20) 61.05 58.73 62.07 63.91 normal-$\uparrow$ 57.17 60.85 $\uparrow$ 53.83 58.83 $\uparrow$ 51.77 56.11 $\uparrow$ 52.39 60.96 $\uparrow$ 54.64 60.27 $\uparrow$ 58.16 57.57 53.52 58.97 $\uparrow$
adult (n=20) 73.84 68.03 52.80 67.10 $\uparrow$ 49.15 66.52 $\uparrow$ 61.57 68.93 $\uparrow$ 70.46 71.19 $\uparrow$ 62.04 71.87 normal-$\uparrow$ 59.07 69.16 $\uparrow$ 65.52 63.33 64.40 70.41 $\uparrow$
drug (n=20) 65.66 63.49 64.40 66.87 normal-$\uparrow$ 62.69 64.21 $\uparrow$ 56.76 64.18 $\uparrow$ 61.04 62.48 $\uparrow$ 52.22 62.62 $\uparrow$ 53.50 61.30 $\uparrow$ 62.34 60.69 55.00 58.89 $\uparrow$
covid (n=40) 64.94 65.10 66.55 67.85 normal-$\uparrow$ 62.54 64.64 $\uparrow$ 57.16 63.41 $\uparrow$ 63.15 60.93 51.66 61.81 $\uparrow$ 57.04 63.13 $\uparrow$ 62.19 60.49 54.51 60.32 $\uparrow$
cutract (n=40) 62.32 63.01 62.49 65.29 normal-$\uparrow$ 61.61 63.01 $\uparrow$ 56.40 62.63 $\uparrow$ 58.18 58.97 $\uparrow$ 51.41 62.50 $\uparrow$ 55.64 62.24 $\uparrow$ 60.93 61.29 $\uparrow$ 51.94 60.24 $\uparrow$
maggic (n=40) 56.85 53.74 53.20 54.71 normal-$\uparrow$ 52.33 52.97 $\uparrow$ 51.85 53.85 $\uparrow$ 52.48 53.45 $\uparrow$ 49.61 52.87 $\uparrow$ 51.24 54.45 $\uparrow$ 53.69 53.37 50.60 52.66 $\uparrow$
seer (n=40) 78.84 72.40 76.02 77.13 $\uparrow$ 75.18 78.26 normal-$\uparrow$ 68.31 73.89 $\uparrow$ 77.94 73.24 51.45 72.32 $\uparrow$ 67.78 75.95 $\uparrow$ 70.32 72.81 $\uparrow$ 68.79 74.48 $\uparrow$
compas (n=40) 61.12 58.23 62.00 63.43 normal-$\uparrow$ 56.73 59.19 $\uparrow$ 52.23 56.84 $\uparrow$ 56.67 57.55 $\uparrow$ 54.52 60.09 $\uparrow$ 54.47 58.92 $\uparrow$ 56.59 56.47 52.62 57.20 $\uparrow$
adult (n=40) 74.13 69.75 49.47 68.90 $\uparrow$ 47.85 66.79 $\uparrow$ 60.55 70.14 $\uparrow$ 64.06 70.56 $\uparrow$ 58.77 71.63 $\uparrow$ 60.26 71.50 $\uparrow$ 69.43 70.58 $\uparrow$ 59.76 72.41 normal-$\uparrow$
drug (n=40) 66.19 64.18 62.86 65.90 normal-$\uparrow$ 60.09 63.43 $\uparrow$ 60.84 65.40 $\uparrow$ 63.05 64.90 $\uparrow$ 54.80 65.25 $\uparrow$ 54.68 62.40 $\uparrow$ 61.21 61.98 $\uparrow$ 53.77 60.23 $\uparrow$
covid (n=100) 63.57 63.17 67.14 68.60 normal-$\uparrow$ 62.40 66.02 $\uparrow$ 61.34 64.63 $\uparrow$ 63.33 66.42 $\uparrow$ 55.86 64.53 $\uparrow$ 55.44 64.91 $\uparrow$ 61.96 63.40 $\uparrow$ 59.75 64.47 $\uparrow$
cutract (n=100) 62.88 62.69 60.96 65.60 $\uparrow$ 62.12 65.95 $\uparrow$ 59.06 66.20 normal-$\uparrow$ 63.38 65.72 $\uparrow$ 52.62 61.73 $\uparrow$ 58.11 65.08 $\uparrow$ 60.15 61.34 $\uparrow$ 58.82 63.76 $\uparrow$
maggic (n=100) 56.67 54.83 53.58 55.70 normal-$\uparrow$ 50.72 53.98 $\uparrow$ 52.29 55.06 $\uparrow$ 53.93 54.61 $\uparrow$ 49.71 53.34 $\uparrow$ 51.82 54.53 $\uparrow$ 54.11 54.34 $\uparrow$ 51.31 53.33 $\uparrow$
seer (n=100) 78.48 73.83 76.50 77.84 normal-$\uparrow$ 74.81 77.80 $\uparrow$ 69.61 76.91 $\uparrow$ 74.07 77.21 $\uparrow$ 59.62 75.40 $\uparrow$ 69.51 76.54 $\uparrow$ 71.04 72.63 $\uparrow$ 66.35 76.44 $\uparrow$
compas (n=100) 61.17 56.61 63.04 63.39 normal-$\uparrow$ 55.57 57.40 $\uparrow$ 55.07 58.95 $\uparrow$ 53.38 55.69 $\uparrow$ 54.61 59.31 $\uparrow$ 54.38 57.92 $\uparrow$ 56.69 57.01 $\uparrow$ 54.34 56.72 $\uparrow$
adult (n=100) 74.22 72.19 46.68 66.58 $\uparrow$ 47.94 67.37 $\uparrow$ 62.42 71.78 $\uparrow$ 68.06 70.77 $\uparrow$ 66.29 74.36 normal-$\uparrow$ 58.28 72.14 $\uparrow$ 71.51 72.68 $\uparrow$ 62.11 72.18 $\uparrow$
drug (n=100) 65.98 64.20 66.18 67.41 normal-$\uparrow$ 60.77 61.21 $\uparrow$ 57.67 63.41 $\uparrow$ 63.26 64.55 $\uparrow$ 54.89 65.72 $\uparrow$ 52.06 63.38 $\uparrow$ 61.21 62.30 $\uparrow$ 55.40 63.32 $\uparrow$
covid (n=200) 64.28 63.45 67.19 68.74 normal-$\uparrow$ 63.25 65.73 $\uparrow$ 61.58 66.27 $\uparrow$ 63.55 66.98 $\uparrow$ 57.20 65.54 $\uparrow$ 55.08 63.77 $\uparrow$ 63.60 63.40 60.01 64.27 $\uparrow$
cutract (n=200) 62.77 63.53 63.36 65.33 $\uparrow$ 63.01 65.47 $\uparrow$ 60.51 65.01 $\uparrow$ 62.55 66.94 normal-$\uparrow$ 57.40 65.14 $\uparrow$ 59.13 65.20 $\uparrow$ 61.66 62.40 $\uparrow$ 58.29 64.85 $\uparrow$
maggic (n=200) 56.72 55.06 52.76 54.41 $\uparrow$ 52.41 54.34 $\uparrow$ 52.47 55.21 normal-$\uparrow$ 52.40 53.39 $\uparrow$ 50.43 53.66 $\uparrow$ 51.58 54.82 $\uparrow$ 54.64 54.36 51.78 54.22 $\uparrow$
seer (n=200) 78.58 74.51 75.20 77.26 $\uparrow$ 75.33 77.41 $\uparrow$ 70.96 77.08 $\uparrow$ 74.51 79.20 normal-$\uparrow$ 71.44 78.66 $\uparrow$ 65.84 76.60 $\uparrow$ 73.11 74.77 $\uparrow$ 66.37 76.36 $\uparrow$
compas (n=200) 60.82 57.20 61.09 62.89 normal-$\uparrow$ 56.15 59.02 $\uparrow$ 54.20 57.78 $\uparrow$ 55.72 58.68 $\uparrow$ 55.83 59.81 $\uparrow$ 52.13 57.25 $\uparrow$ 56.55 57.34 $\uparrow$ 54.07 58.66 $\uparrow$
adult (n=200) 73.61 72.26 44.41 67.99 $\uparrow$ 48.46 67.24 $\uparrow$ 66.56 74.12 $\uparrow$ 71.61 74.89 $\uparrow$ 69.91 76.52 normal-$\uparrow$ 60.70 74.05 $\uparrow$ 73.11 74.08 $\uparrow$ 63.10 73.82 $\uparrow$
drug (n=200) 65.85 67.04 66.29 66.57 $\uparrow$ 61.88 64.34 $\uparrow$ 59.10 66.06 $\uparrow$ 63.59 63.40 59.11 69.04 normal-$\uparrow$ 53.78 64.55 $\uparrow$ 64.56 65.35 $\uparrow$ 56.82 63.02 $\uparrow$

### v3

@@ C.4 Comparison to random noise baseline
covid (n=40) 77.59 71.57 75.23 75.36 $\uparrow$ 71.55 73.37 $\uparrow$ 66.17 70.85 $\uparrow$ 72.21 72.68 $\uparrow$ 56.03 70.80 $\uparrow$ 67.39 71.82 $\uparrow$ 70.02 69.96 61.19 69.77 $\uparrow$
cutract (n=40) 73.14 69.50 71.26 73.57 $\uparrow$ 69.66 71.34 $\uparrow$ 62.52 68.69 $\uparrow$ 64.50 68.05 $\uparrow$ 51.23 68.23 $\uparrow$ 60.74 68.91 $\uparrow$ 69.42 69.19 59.34 67.80 $\uparrow$
maggic (n=40) 71.38 58.76 60.27 61.88 $\uparrow$ 57.87 60.29 $\uparrow$ 55.87 58.47 $\uparrow$ 55.38 58.03 $\uparrow$ 48.50 57.02 $\uparrow$ 54.53 59.01 $\uparrow$ 58.12 58.29 $\uparrow$ 56.19 58.33 $\uparrow$
seer (n=40) 91.80 84.51 86.63 87.40 $\uparrow$ 86.27 86.80 $\uparrow$ 82.48 86.55 $\uparrow$ 81.62 81.58 52.80 83.03 $\uparrow$ 80.04 85.84 $\uparrow$ 82.53 83.62 $\uparrow$ 79.35 85.81 $\uparrow$
compas (n=40) 69.17 62.90 69.24 68.90 61.74 63.21 $\uparrow$ 56.63 62.55 $\uparrow$ 59.07 59.38 $\uparrow$ 58.01 64.95 $\uparrow$ 58.44 64.44 $\uparrow$ 61.81 61.89 $\uparrow$ 56.60 62.26 $\uparrow$
adult (n=40) 88.64 82.16 48.51 76.59 $\uparrow$ 45.92 76.92 $\uparrow$ 73.38 82.53 $\uparrow$ 74.03 79.84 $\uparrow$ 66.60 80.98 $\uparrow$ 69.47 81.56 $\uparrow$ 80.01 80.51 $\uparrow$ 74.20 82.53 $\uparrow$
drug (n=40) 81.06 72.72 77.03 78.85 $\uparrow$ 76.37 76.48 $\uparrow$ 70.99 74.48 $\uparrow$ 74.65 76.53 $\uparrow$ 64.52 75.09 $\uparrow$ 62.58 73.14 $\uparrow$ 71.80 72.15 $\uparrow$ 67.61 74.04 $\uparrow$
covid (n=100) 77.59 73.16 75.55 76.56 $\uparrow$ 72.69 74.36 $\uparrow$ 70.54 73.99 $\uparrow$ 72.99 74.50 $\uparrow$ 63.40 73.98 $\uparrow$ 63.50 73.79 $\uparrow$ 72.64 73.35 $\uparrow$ 70.45 73.76 $\uparrow$
cutract (n=100) 73.14 70.49 72.29 74.01 $\uparrow$ 71.51 73.48 $\uparrow$ 68.64 73.63 $\uparrow$ 71.32 74.02 $\uparrow$ 54.74 69.69 $\uparrow$ 66.97 72.98 $\uparrow$ 67.71 69.42 $\uparrow$ 66.44 72.32 $\uparrow$
maggic (n=100) 71.38 60.55 60.04 62.57 $\uparrow$ 56.98 60.49 $\uparrow$ 58.21 60.42 $\uparrow$ 58.11 59.42 $\uparrow$ 49.71 58.77 $\uparrow$ 55.36 60.58 $\uparrow$ 60.02 60.15 $\uparrow$ 56.47 60.28 $\uparrow$
seer (n=100) 91.80 85.52 86.81 87.64 $\uparrow$ 86.39 87.21 $\uparrow$ 83.74 87.29 $\uparrow$ 84.66 86.85 $\uparrow$ 68.20 85.49 $\uparrow$ 82.37 86.80 $\uparrow$ 83.18 83.87 $\uparrow$ 80.45 86.49 $\uparrow$
compas (n=100) 69.17 62.71 69.37 69.37 59.83 62.80 $\uparrow$ 60.56 64.90 $\uparrow$ 58.20 62.00 $\uparrow$ 58.51 64.60 $\uparrow$ 59.76 64.68 $\uparrow$ 61.33 62.03 $\uparrow$ 60.37 63.42 $\uparrow$
adult (n=100) 88.64 82.78 44.60 76.75 $\uparrow$ 46.34 76.83 $\uparrow$ 74.97 82.89 $\uparrow$ 77.38 81.42 $\uparrow$ 78.84 83.96 $\uparrow$ 68.31 82.87 $\uparrow$ 82.58 83.24 $\uparrow$ 76.15 82.47 $\uparrow$
drug (n=100) 81.06 75.49 79.12 79.81 $\uparrow$ 72.52 74.78 $\uparrow$ 68.89 75.91 $\uparrow$ 76.88 78.33 $\uparrow$ 66.38 77.02 $\uparrow$ 62.27 75.13 $\uparrow$ 74.40 74.49 $\uparrow$ 71.02 76.85 $\uparrow$
covid (n=200) 77.59 73.06 75.12 76.64 $\uparrow$ 71.74 74.66 $\uparrow$ 72.57 75.50 $\uparrow$ 73.56 75.88 $\uparrow$ 66.02 74.51 $\uparrow$ 63.29 74.62 $\uparrow$ 73.34 73.89 $\uparrow$ 68.14 74.21 $\uparrow$
cutract (n=200) 73.14 71.68 73.43 75.14 $\uparrow$ 71.76 74.31 $\uparrow$ 69.35 73.80 $\uparrow$ 72.76 75.33 $\uparrow$ 65.83 73.43 $\uparrow$ 68.05 74.27 $\uparrow$ 70.06 71.35 $\uparrow$ 66.65 73.34 $\uparrow$
maggic (n=200) 71.38 62.40 59.46 63.15 $\uparrow$ 58.05 61.68 $\uparrow$ 59.04 62.58 $\uparrow$ 57.63 59.23 $\uparrow$ 50.63 60.09 $\uparrow$ 55.55 61.36 $\uparrow$ 61.73 61.88 $\uparrow$ 57.09 61.15 $\uparrow$
seer (n=200) 91.80 85.87 86.75 88.08 $\uparrow$ 85.74 86.97 $\uparrow$ 83.97 87.64 $\uparrow$ 86.21 88.46 $\uparrow$ 83.21 87.65 $\uparrow$ 76.91 87.16 $\uparrow$ 84.70 85.59 $\uparrow$ 81.45 87.46 $\uparrow$
compas (n=200) 69.17 63.14 68.80 69.54 $\uparrow$ 61.62 64.59 $\uparrow$ 60.65 65.67 $\uparrow$ 60.97 65.12 $\uparrow$ 60.66 65.67 $\uparrow$ 56.49 64.34 $\uparrow$ 61.68 62.93 $\uparrow$ 60.32 64.18 $\uparrow$
adult (n=200) 88.64 84.12 40.38 77.39 $\uparrow$ 47.40 79.40 $\uparrow$ 80.00 84.61 $\uparrow$ 83.73 85.73 $\uparrow$ 81.92 85.76 $\uparrow$ 74.90 85.13 $\uparrow$ 84.78 84.45 75.54 84.39 $\uparrow$
drug (n=200) 81.06 78.41 78.25 79.09 $\uparrow$ 73.31 76.76 $\uparrow$ 72.82 79.00 $\uparrow$ 72.62 74.89 $\uparrow$ 74.30 79.75 $\uparrow$ 63.91 77.60 $\uparrow$ 77.64 77.92 $\uparrow$ 70.16 77.89 $\uparrow$
@@
covid (n=20) 63.46 61.92 67.52 67.67 $\uparrow$ 61.07 64.89 $\uparrow$ 52.94 56.62 $\uparrow$ 58.95 59.19 $\uparrow$ 51.83 58.73 $\uparrow$ 54.37 61.53 $\uparrow$ 59.56 59.01 56.46 59.49 $\uparrow$
cutract (n=20) 62.49 64.62 62.40 66.03 $\uparrow$ 62.10 65.15 $\uparrow$ 55.67 61.73 $\uparrow$ 57.91 58.97 $\uparrow$ 51.89 62.76 $\uparrow$ 57.12 64.85 $\uparrow$ 62.85 62.38 60.25 62.34 $\uparrow$
maggic (n=20) 56.53 54.80 54.67 55.03 $\uparrow$ 52.54 53.54 $\uparrow$ 50.90 52.08 $\uparrow$ 52.39 52.70 $\uparrow$ 49.70 53.08 $\uparrow$ 51.52 53.23 $\uparrow$ 52.55 52.44 50.91 51.52 $\uparrow$
seer (n=20) 78.86 71.09 76.17 77.17 $\uparrow$ 73.92 75.58 $\uparrow$ 64.18 70.69 $\uparrow$ 74.78 74.72 46.60 68.79 $\uparrow$ 65.77 72.25 $\uparrow$ 70.54 71.19 $\uparrow$ 64.47 70.11 $\uparrow$
compas (n=20) 61.05 58.73 62.07 63.91 $\uparrow$ 57.17 60.85 $\uparrow$ 53.83 58.83 $\uparrow$ 51.77 56.11 $\uparrow$ 52.39 60.96 $\uparrow$ 54.64 60.27 $\uparrow$ 58.16 57.57 53.52 58.97 $\uparrow$
adult (n=20) 73.84 68.03 52.80 67.10 $\uparrow$ 49.15 66.52 $\uparrow$ 61.57 68.93 $\uparrow$ 70.46 71.19 $\uparrow$ 62.04 71.87 $\uparrow$ 59.07 69.16 $\uparrow$ 65.52 63.33 64.40 70.41 $\uparrow$
drug (n=20) 65.66 63.49 64.40 66.87 $\uparrow$ 62.69 64.21 $\uparrow$ 56.76 64.18 $\uparrow$ 61.04 62.48 $\uparrow$ 52.22 62.62 $\uparrow$ 53.50 61.30 $\uparrow$ 62.34 60.69 55.00 58.89 $\uparrow$
covid (n=40) 63.46 65.10 66.55 67.85 $\uparrow$ 62.54 64.64 $\uparrow$ 57.16 63.41 $\uparrow$ 63.15 60.93 51.66 61.81 $\uparrow$ 57.04 63.13 $\uparrow$ 62.19 60.49 54.51 60.32 $\uparrow$
cutract (n=40) 62.49 63.01 62.49 65.29 $\uparrow$ 61.61 63.01 $\uparrow$ 56.40 62.63 $\uparrow$ 58.18 58.97 $\uparrow$ 51.41 62.50 $\uparrow$ 55.64 62.24 $\uparrow$ 60.93 61.29 $\uparrow$ 51.94 60.24 $\uparrow$
maggic (n=40) 56.53 53.74 53.20 54.71 $\uparrow$ 52.33 52.97 $\uparrow$ 51.85 53.85 $\uparrow$ 52.48 53.45 $\uparrow$ 49.61 52.87 $\uparrow$ 51.24 54.45 $\uparrow$ 53.69 53.37 50.60 52.66 $\uparrow$
seer (n=40) 78.86 72.40 76.02 77.13 $\uparrow$ 75.18 78.26 $\uparrow$ 68.31 73.89 $\uparrow$ 77.94 73.24 51.45 72.32 $\uparrow$ 67.78 75.95 $\uparrow$ 70.32 72.81 $\uparrow$ 68.79 74.48 $\uparrow$
compas (n=40) 61.05 58.23 62.00 63.43 $\uparrow$ 56.73 59.19 $\uparrow$ 52.23 56.84 $\uparrow$ 56.67 57.55 $\uparrow$ 54.52 60.09 $\uparrow$ 54.47 58.92 $\uparrow$ 56.59 56.47 52.62 57.20 $\uparrow$
adult (n=40) 74.84 69.75 49.47 68.90 $\uparrow$ 47.85 66.79 $\uparrow$ 60.55 70.14 $\uparrow$ 64.06 70.56 $\uparrow$ 58.77 71.63 $\uparrow$ 60.26 71.50 $\uparrow$ 69.43 70.58 $\uparrow$ 59.76 72.41 $\uparrow$
drug (n=40) 65.66 64.18 62.86 65.90 $\uparrow$ 60.09 63.43 $\uparrow$ 60.84 65.40 $\uparrow$ 63.05 64.90 $\uparrow$ 54.80 65.25 $\uparrow$ 54.68 62.40 $\uparrow$ 61.21 61.98 $\uparrow$ 53.77 60.23 $\uparrow$
covid (n=100) 63.46 63.17 67.14 68.60 $\uparrow$ 62.40 66.02 $\uparrow$ 61.34 64.63 $\uparrow$ 63.33 66.42 $\uparrow$ 55.86 64.53 $\uparrow$ 55.44 64.91 $\uparrow$ 61.96 63.40 $\uparrow$ 59.75 64.47 $\uparrow$
cutract (n=100) 62.49 62.69 60.96 65.60 $\uparrow$ 62.12 65.95 $\uparrow$ 59.06 66.20 $\uparrow$ 63.38 65.72 $\uparrow$ 52.62 61.73 $\uparrow$ 58.11 65.08 $\uparrow$ 60.15 61.34 $\uparrow$ 58.82 63.76 $\uparrow$
maggic (n=100) 56.53 54.83 53.58 55.70 $\uparrow$ 50.72 53.98 $\uparrow$ 52.29 55.06 $\uparrow$ 53.93 54.61 $\uparrow$ 49.71 53.34 $\uparrow$ 51.82 54.53 $\uparrow$ 54.11 54.34 $\uparrow$ 51.31 53.33 $\uparrow$
seer (n=100) 78.86 73.83 76.50 77.84 $\uparrow$ 74.81 77.80 $\uparrow$ 69.61 76.91 $\uparrow$ 74.07 77.21 $\uparrow$ 59.62 75.40 $\uparrow$ 69.51 76.54 $\uparrow$ 71.04 72.63 $\uparrow$ 66.35 76.44 $\uparrow$
compas (n=100) 61.05 56.61 63.04 63.39 $\uparrow$ 55.57 57.40 $\uparrow$ 55.07 58.95 $\uparrow$ 53.38 55.69 $\uparrow$ 54.61 59.31 $\uparrow$ 54.38 57.92 $\uparrow$ 56.69 57.01 $\uparrow$ 54.34 56.72 $\uparrow$
adult (n=100) 73.84 72.19 46.68 66.58 $\uparrow$ 47.94 67.37 $\uparrow$ 62.42 71.78 $\uparrow$ 68.06 70.77 $\uparrow$ 66.29 74.36 $\uparrow$ 58.28 72.14 $\uparrow$ 71.51 72.68 $\uparrow$ 62.11 72.18 $\uparrow$
drug (n=100) 65.66 64.20 66.18 67.41 $\uparrow$ 60.77 61.21 $\uparrow$ 57.67 63.41 $\uparrow$ 63.26 64.55 $\uparrow$ 54.89 65.72 $\uparrow$ 52.06 63.38 $\uparrow$ 61.21 62.30 $\uparrow$ 55.40 63.32 $\uparrow$
covid (n=200) 63.46 63.45 67.19 68.74 $\uparrow$ 63.25 65.73 $\uparrow$ 61.58 66.27 $\uparrow$ 63.55 66.98 $\uparrow$ 57.20 65.54 $\uparrow$ 55.08 63.77 $\uparrow$ 63.60 63.40 60.01 64.27 $\uparrow$
cutract (n=200) 62.49 63.53 63.36 65.33 $\uparrow$ 63.01 65.47 $\uparrow$ 60.51 65.01 $\uparrow$ 62.55 66.94 $\uparrow$ 57.40 65.14 $\uparrow$ 59.13 65.20 $\uparrow$ 61.66 62.40 $\uparrow$ 58.29 64.85 $\uparrow$
maggic (n=200) 56.53 55.06 52.76 54.41 $\uparrow$ 52.41 54.34 $\uparrow$ 52.47 55.21 $\uparrow$ 52.40 53.39 $\uparrow$ 50.43 53.66 $\uparrow$ 51.58 54.82 $\uparrow$ 54.64 54.36 51.78 54.22 $\uparrow$
seer (n=200) 78.86 74.51 75.20 77.26 $\uparrow$ 75.33 77.41 $\uparrow$ 70.96 77.08 $\uparrow$ 74.51 79.20 $\uparrow$ 71.44 78.66 $\uparrow$ 65.84 76.60 $\uparrow$ 73.11 74.77 $\uparrow$ 66.37 76.36 $\uparrow$
compas (n=200) 61.05 57.20 61.09 62.89 $\uparrow$ 56.15 59.02 $\uparrow$ 54.20 57.78 $\uparrow$ 55.72 58.68 $\uparrow$ 55.83 59.81 $\uparrow$ 52.13 57.25 $\uparrow$ 56.55 57.34 $\uparrow$ 54.07 58.66 $\uparrow$
adult (n=200) 73.84 72.26 44.41 67.99 $\uparrow$ 48.46 67.24 $\uparrow$ 66.56 74.12 $\uparrow$ 71.61 74.89 $\uparrow$ 69.91 76.52 $\uparrow$ 60.70 74.05 $\uparrow$ 73.11 74.08 $\uparrow$ 63.10 73.82 $\uparrow$
drug (n=200) 65.66 67.04 66.29 66.57 $\uparrow$ 61.88 64.34 $\uparrow$ 59.10 66.06 $\uparrow$ 63.59 63.40 59.11 69.04 $\uparrow$ 53.78 64.55 $\uparrow$ 64.56 65.35 $\uparrow$ 56.82 63.02 $\uparrow$